Mrs Chopra is 5 time as old as her son now,In 12 years' time, Mrs Chopra will be 3 time as old as her son. What is Mrs Chopra's age now?
Answer by Cromlix(4381)

hi there,
Mrs Chopra is 5 time as old as her son now
MrsC = 5S
In 12 years' time, Mrs Chopra will be 3 time
as old as her son.
MrsC + 12 = 3(S + 12)
Replace MrsC with 5S (Top equation)
5S + 12 + 3(S + 12)
5S + 12 = 3S + 36
Collect like terms:
5S - 3S = 36 - 12
2S = 24
S = 12
Mrs Chopra is 60 years old now
Son is 12 years old now.
Hope this helps :-)